college algebra - Fifty liters of pure maple syrup that costs $9.50 per liter are mixed with imitation maple syrup that costs $ 4.00 per liter. How much imitation maple syrup is needed to make a mixture that costs $5.00 per liter?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

Fifty liters of pure maple syrup that costs $9.50 per liter are mixed with imitation maple syrup that costs $ 4.00 per liter. How much imitation maple syrup is needed to make a mixture that costs $5.00 per liter?
-----------------
Equation:
value + value = value
50*9.5 + 4x = 5(50+x)
----
Multiply thru by 10 to get;
50*95 + 40x = 50*50 + 50x
10x = 45*50
x = 9 liters (amt of $4 syrup needed)
=============================-----------
Cheers,
Stan H.

Answer by nerdybill(7384)

Fifty liters of pure maple syrup that costs $9.50 per liter are mixed with imitation maple syrup that costs $ 4.00 per liter. How much imitation maple syrup is needed to make a mixture that costs $5.00 per liter?
.
Let x = amount (liters) of imitation maple
then
50(9.50) + x(4) = (50+x)5
475 + 4x = 250+5x
475 = 250+x
200 liters = x